OpenStudy (jiteshmeghwal9):

Simplify:- \[{(2x+5)(4x^2-20x+25)-(2x-5)(4x^2+20x+25)}\] I got \[-40x^2+250\]. Is it right ??

OpenStudy (jiteshmeghwal9):

If no then what is the answer?? give full solution plz :)

OpenStudy (jiteshmeghwal9):

@him1618 @lgbasallote plz help :)

OpenStudy (compassionate):

If you want me too, I can sit down and work this homogeneous problem out.

OpenStudy (jiteshmeghwal9):

sure :)

OpenStudy (lgbasallote):

if you factor it out you get (2x+5)(2x-5)^2 - (2x-5)(2x+5)^2 then if you factor out further.. (2x+5)(2x-5)(2x - 5 - (2x+5) ) right?

OpenStudy (lgbasallote):

so you should get (4x^2 - 25)(-10) then expand

OpenStudy (lgbasallote):

so you're right

OpenStudy (jiteshmeghwal9):

but the answer given in my book is \(-40x^2+40x+150\) :(

OpenStudy (jiteshmeghwal9):

is the answer given in my book is wrong or i m wrong ??

OpenStudy (lgbasallote):

the question is wrong...?

OpenStudy (jiteshmeghwal9):

no, i m talking about the answer

OpenStudy (jiteshmeghwal9):

@lgbasallote u there ??

OpenStudy (anonymous):

In any case an x term cannot be left out. So yes, you are right. Book s wrong.

OpenStudy (anonymous):

\[(2x+5)(2x-5)^2-(2x-5)(2x+5)^2=(2x+5)(2x-5)[(2x-5) - (2x+5)]=4x^2-25(-10)=250-40x^2\] do u get it

OpenStudy (anonymous):

i think the answer has gone out of page can u see it its...250-40x^2

OpenStudy (jiteshmeghwal9):

no ! the exercise is 9.6 & the first question of fourth question

OpenStudy (jiteshmeghwal9):

Thanx @lgbasallote sir :)